# Intelligent Surfaces Empowered Wireless Network: Recent Advances and The Road to 6G

### 15 Jan 2024 v2

@@ II-D1 IS-enhanced Wireless Sensing
Promising new system architectures are essential to overcome the sensing range bottleneck for wireless sensing systems. IS passive sensing is shown in Fig. 13(a), where the sensing performance is improved by establishing a virtual LoS link from the BS to the IS and then to the target [157]. However, the practical performance of IS passive sensing is limited by the severe path loss of the cascaded echo link. To address this issue, IS semi-passive sensing was proposed in [158] (see Fig. 13(b)), where dedicated low-cost active sensors are installed on the IS (i.e., semi-passive IS), allowing it not only to reflect signals from the BS but also to directly receive echo signals from the target for sensing. As such, IS semi-passive sensing incurs lower path loss compared with IS passive sensing. Nevertheless, if the BS is far from the IS, the sensing accuracy of IS semi-passive sensing remains limited. To overcome this issue, a new type of IS-aided sensing architecture, called IS active sensing, has recently been proposed [159] (see Fig. 13(c)). Specifically, in IS active sensing, the IS controller, typically responsible for reflection control and information exchange only, also acts as a transmitter to send probing signals for sensing. This allows the IS to autonomously send (via the IS controller) and receive (via sensors) sensing signals for target localization, thus avoiding the dependence on BSs for sensing signal transmission/reception. Although IS active sensing incurs higher hardware and energy cost than IS passive/semi-passive sensing, it achieves superior sensing performance due to the short signal propagation distance and also helps reduce the BS transmit power for sensing.
@@
Besides enhancing sensing accuracy, target-mounted IS has also been proposed to improve wireless sensing security. Specifically, the secure sensing aims to enhance sensing performance of the legitimate radar station (LRS) while preventing the target detection by any unauthorized radar station (URS). To achieve this goal, the authors in [160] proposed a general secure sensing protocol with target-mounted semi-passive IS, where each radar coherent-processing interval is divided into two phases. In the first phase, the IS sensors estimate the LRS/URS channel and waveform parameters with all IS REs switched off. After that, in the second phase, based on the estimated parameters, IS reflection is designed to simultaneously boost the received signal at the LRS receiver and suppress that at the URS receiver. Benefiting from the flexible passive beamforming gain, such a new design with target-mounted IS can offer significant secure sensing performance gains as compared to the conventional sensing without target-mounted IS.
@@ II-D2 IS-aided ISAC
How to design IS-assisted ISAC systems to achieve optimal sensing and communication performance without significantly changing existing protocols in cellular systems is a key challenge. To overcome this challenge, the authors in [164] proposed to exploit cooperative ISs’ passive beam searching and the communication signal from the BS (e.g., pilot signal sent for channel estimation) for cellular UAV target sensing. Accordingly, a general transmission protocol for sensing is illustrated in Fig. 14, where each channel coherence interval is divided into three phases. In the first phase, background channel is estimated without IS reflection. After that, in the second phase, the angle parameters between IS 1 and UAV are estimated with IS 2 OFF. Finally, in the third phase, the angle parameters between IS 2 and UAV are estimated with IS 1’s beam fixed as the best one obtained in the second phase. Note that by properly selecting the idle BS (i.e., without communicating with any user) and its sensing resource block, the communication interference from other active BSs (i.e., communicating with a user) to the sensing signal at idle BS can be ignored. The authors demonstrated that this cellular ISAC scheme not only requires no modifications to the existing BS beamforming design and transmission protocol, but also does not impact the communication performance of users in the existing network, thus achieving a superior performance tradeoff between sensing and communication.
The IS is also particularly advantageous in mmWave ISAC to compensate for the high path loss, thus greatly improving the sensing coverage and communications performance. In mmWave communication, beam scanning/training is essential[165], which can be exploited for simultaneous target sensing. Motivated by this, in [166], the authors considered a mmWave ISAC system where a multiple-antenna BS attempts to communicate with a single-antenna user, and also to detect the angle of a sensing target from the semi-passive IS. They proposed a two-phase simultaneous beam training and sensing (STAS) protocol to enhance local communication and sensing performance at the same time via IS passive beam scanning. Specifically, in the first phase, the ISAC system finds the best beam for communication user and concurrently detects the target based on its echo signal received by IS sensors. Next, in the second phase, data are sent to the communication user with the best IS beam found in the first phase. Due to the effective utilization of the downlink beam training signal for simultaneous sensing, the STAS protocol outperforms the benchmark protocol with orthogonal beam training and sensing.
@@
TABLE V: A Summary of Representative Works on IS-Aided Sensing/ISAC Systems.
@@
In Table IV, we summarize the up-to-date research works on exploring IS-aided sensing and ISAC. To motivate future research on IS-aided ISAC, several important directions are briefly discussed as follows which still need in-depth investigation. First, it is necessary to explore the connection between the sensing and communication channels to enhance ISAC performance. Furthermore, new IS architectures, such as target-mounted IS[162], STAR-RIS [173], and active IS[174] (to be discussed in the next section in detail), open new opportunities for ISAC performance improvement.
@@ III-A1 Active IS
The communication performance of passive-IS-assisted systems is practically constrained by the severe product-distance path-loss in signal reflections, which can be compensated for by equipping ISs with a massive number of passive reflecting elements or reduced by deploying the passive ISs closer to the transceivers. However, these methods may not be applicable in practice, due to the high design complexity and practical site constraints. To address this issue, an efficient approach is by deploying a new type of IS for assisting wireless communication, called active IS, which enables the signal reflection with power amplification for effectively compensating for the product-distance path-loss [175, 176]. Generally speaking, active IS can be implemented by two types of hardware architectures as illustrated in Fig. 16, namely, 1) the cascaded amplifying and phase-shifting (APS) circuits (see Fig. 16 (a)) [175] and 2) load modulation circuits (e.g., tunnel diodes) (see Fig. 16 (b)) [176]. With the power amplification capability, active IS can be used for not only improving the communication performance, but also enhancing the performance of other wireless applications, such as wireless power transfer [177] and physical layer security [178]. In this subsection, we provide an overview of recent research results on active IS from the communication design perspective, including active-IS reflection optimization, channel estimation, and deployment.
@@
Last, for the active-IS deployment design, several factors need to be considered, such as the physical environment, number of ISs and their REs, uplink or downlink communication, amplification power budget, etc. Specifically, for the single-active-IS aided communication system, it was shown in [174] that the active-IS should be deployed closer to the receiver to compensate for the amplification noise in the downlink/uplink. It was revealed in [184] that given the fixed per-element amplification power, the received SNR at the user increases asymptotically with the square of the number of REs, while the double-active-IS aided wireless system only achieves a linear capacity scaling order with the number of active REs under the total amplification power constraint, due to the noise amplification over the double-reflection link.
@@ III-A2 Integrated Active/Passive IS
In future work, there are several key design issues worthy of investigation. For example, the existing works on active-IS have usually assumed ideal reflection coefficients with continuous and independent phase-shift and amplitude control, which may not be implementable in practice and thus calls for efficient algorithms to handle their practical constraints. In addition, for MAMP-IS aided systems, it is interesting to study how to optimize the density of active and passive ISs in a multi-cell network for maximizing the network throughput by using e.g., the stochastic geometry theory [190].
@@ III-B Omnidirectional IS
Our discussions so far have mainly focused on reflection-based ISs, passive or active. However, reflection-based ISs have one fundamental limitation after they are deployed in practice: they can only reflect incident waves from a restricted angular range, i.e., only users and the BS that are located in the same reflection half-space of the IS can benefit from the signal reflection by the IS. To overcome this restriction, a new concept of omnidirectional IS has been proposed [191]. The omnidirectional IS can enable a dual function of signal reflection and refraction simultaneously, i.e., some power of the incident signal is reflected to serve users located in the same reflection half-space of the IS, and the rest of power penetrates the IS to serve users located in the refraction half-space of the IS. As a result, a doubled ${}^{\circ}$ service coverage is ensured compared to reflection-based IS. This type of surfaces is also referred to as STAR-RIS as aforementioned [192]. In this subsection, we first introduce the working principles of the omnidirectional IS and its new characteristics. Then, two hardware implementations of the omnidirectional IS are presented. Finally, emerging applications of the omnidirectional IS are discussed.
@@ III-B3 Hardware Implementation
It is worth noting that some existing designs utilized coupling among elements to realize simultaneous reflection and refraction [19]. In other words, the incident signals on an element will also be reflected or refracted from the neighboring elements. As a result, the phase shift matrix $\bm{\Phi}$ is not diagonal any more for either reflection or refraction, and this type of surface is also referred to as BD-RIS as mentioned in Section I-B. Such a coupling can provide another dimension to optimize, i.e., the equivalent amplitude of the coefficient for each element can be adjusted, leading to additional performance improvement. However, the coupling poses an increasing complexity on the hardware designs of the omnidirectional IS. Therefore, it is necessary to strike a trade-off between the performance and complexity under such a framework.
@@ III-C3 Hardware Implementation
The primary challenge in the hardware implementation of holographic ISs is to design tunable radiation elements which are capable of optimizing the current distribution on the surface. Specifically, the tunability of radiation elements enables the construction of any holographic pattern on the surface, such that dynamic beamforming can be achieved. In the following we will list the state-of-the-art tuning mechanisms for the realization of holographic ISs, which rely on PIN diodes, liquid crystal, and photosensitive devices, respectively.
@@
  • PIN diodes: A radiating element capable of controlling the radiation amplitude of the EM waves can be achieved through the integration of a complementary electric-LC (cELC) resonator, which incorporates PIN diodes for amplitude modulation [212]. The cELC resonator, as depicted in Fig. 21, is a micro-strip line etched with a circle of annular slot to form a metal patch combined with a closed loop. The operational states of PIN diodes are actively controlled by the application of distinct bias voltages. Such control holds significant influence over the mutual inductance properties of the cELC resonator, consequently dictating the radiation power of the EM wave. Moreover, it was verified that a holographic IS loaded with PIN diodes capable of binary discrete amplitude control is enough for beampattern gain maximization [212].
  • Liquid crystal: Although the usage of PIN diodes is enough for beampattern gain maximization, the nonlinearities of PIN diodes in high-frequency bands will deteriorate the performance of holographic ISs. Due to the linear properties of liquid crystal in high-frequency bands, the above limitation of PIN diodes can be overcome. The liquid crystal-based holographic ISs were proposed in [213], where the liquid crystal is filled in cavity boxes between slots and radiating patches. The permittivity of liquid crystal is changed with the applied bias voltages, such that the EM wave can be regulated to achieve beamforming.
  • Photosensitive devices: In addition to the above electrical control methods, the tunability of radiation elements can also be realized by optical control. Specifically, different EM responses of each radiation element are controlled by properly illuminating the corresponding photosensitive semiconductors. Such an optical control method is capable of achieving a rapid tuning speed while safeguarding signals against EM interference. Moreover, when integrated with optical fiber, it becomes feasible to deploy holographic ISs and signal processing units at separate locations for long-distance deployments [214].
@@ III-C4 Potential Applications
  • Low-Earth-orbit (LEO) satellite communications: LEO satellite communications are one promising solution to deliver high-capacity backhaul or data relay services for terrestrial networks. Nonetheless, the high mobility of LEO satellites and the severe path loss pose challenges on antenna technologies in terms of precise beam steering and high antenna gain [215]. Traditional antennas integrated with user terminals such as dish antennas and phased arrays either require heavy mechanics or costly phase shifters, rendering their adoption in practical systems economically unviable. As a new paradigm to achieve beamforming with low hardware cost and low power consumption, holographic ISs can be adopted in integrated terrestrial-satellite networks to overcome the above limitations. In particular, Kymeta together with Microsoft has developed a commercial holographic IS for satellite communications in the event of natural disasters.
  • Wireless Simultaneous Localization and Mapping (SLAM): Wireless SLAM, as a technique for user positioning and mapping in uncharted environments, holds promise for enhancing location-based services. This method utilizes antennas to estimate both the time of arrival and the angle of arrival by analyzing the amplitudes of received multi-path components [216]. Consequently, the precision of the SLAM system is determined by the directive gain of the antennas. Given that holographic ISs consist of numerous elements, offering enhanced beam-steering capabilities and high directive gain, they present a viable option to be applied in wireless SLAM systems. Moreover, holographic ISs also show the potential in ISAC [217].

### 24 Mar 2024 v3

@@ II-D1 IS-enhanced Wireless Sensing
Promising new system architectures are essential to overcome the sensing range bottleneck for wireless sensing systems. IS passive sensing is shown in Fig. 13(a), where the sensing performance is improved by establishing a virtual LoS link from the BS to the IS and then to the target [239, 240]. However, the practical performance of IS passive sensing is limited by the severe path loss of the cascaded echo link. To address this issue, IS semi-passive sensing was proposed in [241] (see Fig. 13(b)), where dedicated low-cost active sensors are installed on the IS (i.e., semi-passive IS), allowing it not only to reflect signals from the BS but also to directly receive echo signals from the target for sensing. As such, IS semi-passive sensing incurs lower path loss compared with IS passive sensing. Nevertheless, if the BS is far from the IS, the sensing accuracy of IS semi-passive sensing remains limited. To overcome this issue, a new type of IS-aided sensing architecture, called IS active sensing, has recently been proposed [242] (see Fig. 13(c)). Specifically, in IS active sensing, the IS controller, typically responsible for reflection control and information exchange only, also acts as a transmitter to send probing signals for sensing. This allows the IS to autonomously send (via the IS controller) and receive (via sensors) sensing signals for target localization, thus avoiding the dependence on BSs for sensing signal transmission/reception. Although IS active sensing incurs higher hardware and energy cost than IS passive/semi-passive sensing, it achieves superior sensing performance due to the short signal propagation distance and also helps reduce the BS transmit power for sensing.
@@
Besides enhancing sensing accuracy, target-mounted IS has also been proposed to improve wireless sensing security. Specifically, the secure sensing aims to enhance sensing performance of the legitimate radar station (LRS) while preventing the target detection by any unauthorized radar station (URS). To achieve this goal, the authors in [243] proposed a general secure sensing protocol with target-mounted semi-passive IS, where each radar coherent-processing interval is divided into two phases. In the first phase, the IS sensors estimate the LRS/URS channel and waveform parameters with all IS REs switched off. After that, in the second phase, based on the estimated parameters, IS reflection is designed to simultaneously boost the received signal at the LRS receiver and suppress that at the URS receiver. Benefiting from the flexible passive beamforming gain, such a new design with target-mounted IS can offer significant secure sensing performance gains as compared to the conventional sensing without target-mounted IS. Moreover, target-mounted IS also gives rise to a novel application in electromagnetic stealth technology for evading radar detection [247, 248], which offers a flexible and adaptable solution to overcome the limitations of traditional stealth materials. In [247, 248], the authors proposed an optimization problem for designing the IRS’s reflection at the target to minimize the sum of received signal power over all adversary radars to achieve electromagnetic stealth. Low-complexity closed-form solutions based on reverse alignment/cancellation and minimum mean square error (MMSE) criteria were proposed in [247] for the single-radar and multi-radar cases, respectively. This new application of target-mounted IS opens up new research avenues in anti-radar detection, providing a promising direction for future exploration in this field.
@@ II-D2 IS-aided ISAC
How to design IS-assisted ISAC systems to achieve optimal sensing and communication performance without significantly changing existing protocols in cellular systems is a key challenge [249, 250, 251]. To overcome this challenge, the authors in [252] proposed to exploit cooperative ISs’ passive beam searching and the communication signal from the BS (e.g., pilot signal sent for channel estimation) for cellular UAV target sensing. The associated sensing protocol is shown in Fig. 14, where each channel coherence interval is divided into three phases. In the first phase, background channel is estimated without IS reflection. After that, in the second phase, the angle parameters between IS 1 and UAV are estimated with IS 2 OFF. Finally, in the third phase, the angle parameters between IS 2 and UAV are estimated with IS 1’s beam fixed as the best one obtained in the second phase. Note that by properly selecting the idle BS (i.e., without communicating with any user) and its sensing resource block, the communication interference from other active BSs (i.e., communicating with a user) to the sensing signal at idle BS can be ignored thanks to the inter-cell interference coordination (ICIC). The authors demonstrated that this cellular ISAC scheme not only requires no modifications to the existing BS beamforming design and transmission protocol, but also does not impact the communication performance of users in the existing network, thus achieving a superior performance tradeoff between sensing and communication.
The IS is also particularly advantageous in mmWave ISAC to compensate for the high path loss, thus greatly improving the sensing coverage and communications performance [253, 254]. In mmWave communication, beam scanning/training is widely applied[255] and can be exploited for simultaneous target sensing. In [256], the authors considered a mmWave ISAC system where a multiple-antenna BS aims to communicate with a single-antenna user, and also to detect the angle of a sensing target from the semi-passive IS. They proposed a two-phase simultaneous beam training and sensing (STAS) protocol to enhance local communication and sensing performance at the same time via IS passive beam scanning. Specifically, in the first phase, the ISAC system finds the best beam for communication user and concurrently detects the target based on its echo signal received by IS sensors. Next, in the second phase, data are sent to the communication user with the best IS beam found in the first phase. Due to the effective utilization of the downlink beam training signal for simultaneous sensing, the STAS protocol outperforms the benchmark protocol with orthogonal beam training and sensing.
@@
TABLE VI: A Summary of Representative Works on IS-Aided Sensing/ISAC Systems.
@@
Sensing [247, 248]
Electromagnetic stealth
Only LoS
Semi-passive IS
Stealth and sensing security
Optimize the IS’s reflection at the target to minimize the received signal power of adversary radars
Reverse cancellation/MMSE and Lagrange multiplier
@@
Summary: In this subsection, we have discussed IS-aided sensing schemes based on four approaches, namely, IS passive sensing, IS semi-passive sensing, IS active sensing, and target-mounted IS sensing. We compared their pros and cons in terms of performance, hardware cost, and implementation complexity. In addition, we discussed IS-aided ISAC for cellular-connected UAV and mmWave systems. In Table IV, we summarize the up-to-date research works on exploring IS-aided sensing and ISAC. To motivate future research on IS-aided ISAC, several important directions are briefly discussed as follows which still need in-depth investigation. First, it is necessary to explore the connection between the IS-aided sensing and communication channels to enhance ISAC performance. Furthermore, new IS architectures, such as target-mounted IS[245], STAR-RIS [263], and active IS[264] (to be discussed in the next section in detail), open new opportunities for ISAC performance improvement.
@@ III-A1 Active IS
The communication performance of passive-IS-assisted systems is practically constrained by the severe product-distance path-loss in signal reflections, which can be compensated for by equipping ISs with a massive number of passive reflecting elements or reduced by deploying the passive ISs closer to the transceivers. However, these methods may not be applicable in practice, due to the high design complexity and practical site constraints. To address this issue, an efficient approach is by deploying a new type of IS for assisting wireless communication, called active IS, which enables the signal reflection with power amplification for effectively compensating for the product-distance path-loss [265, 266, 49]. Generally speaking, active IS can be implemented by two types of hardware architectures as illustrated in Fig. 16, namely, 1) the cascaded amplifying and phase-shifting (APS) circuits (see Fig. 16 (a)) [265] and 2) load modulation circuits (e.g., tunnel diodes) (see Fig. 16 (b)) [266]. With the power amplification capability, active IS can be used for not only improving the communication performance, but also enhancing the performance of other wireless applications, such as wireless power transfer, physical layer security [267, 268, 269, 270, 271, 272, 273, 274]. In this subsection, we provide an overview of recent research results on active IS from the communication design perspective, including active-IS reflection optimization, channel estimation, and deployment.
@@
Last, for the active-IS deployment design, several factors need to be considered, such as the physical environment, number of ISs and their REs, uplink or downlink communication, amplification power budget, etc. Specifically, for the single-active-IS aided communication system, it was shown in [264] that the active-IS should be deployed closer to the receiver to compensate for the amplification noise in the downlink/uplink. It was revealed in [280] that given the fixed per-element amplification power, the received SNR at the user increases asymptotically with the square of the number of REs, while the double-active-IS aided wireless system only achieves a linear capacity scaling order with the number of active REs under the total amplification power constraint, due to the noise amplification over the double-reflection link [281].
@@ III-A2 Integrated Active/Passive IS
In future work, there are several key design issues worthy of investigation. For example, the existing works on active-IS have usually assumed ideal reflection coefficients with continuous and independent phase-shift and amplitude control, which may not be implementable in practice and thus calls for efficient algorithms to handle their practical constraints. In addition, for MAMP-IS aided systems, it is interesting to study how to optimize the density of active and passive ISs in a multi-cell network for maximizing the network throughput by using e.g., the stochastic geometry theory [287]. Furthermore, a research direction is the design of globally passive ISs based on the optimization of surface waves, since no power amplifiers are required but more advanced optimization schemes are needed [288].
@@ III-B Omnidirectional IS
Our discussions so far have mainly focused on reflection-based ISs, passive or active. However, reflection-based ISs have one fundamental limitation after they are deployed in practice: they can only reflect incident waves from a restricted angular range, i.e., only users and the BS that are located in the same reflection half-space of the IS can benefit from the signal reflection by the IS. To overcome this restriction, a new concept of omnidirectional IS has been proposed [289]. The omnidirectional IS can enable a dual function of signal reflection and refraction simultaneously, i.e., some power of the incident signal is reflected to serve users located in the same reflection half-space of the IS, and the rest of power penetrates the IS to serve users located in the refraction half-space of the IS [290, 291]. As a result, a doubled ${}^{\circ}$ service coverage is ensured compared to reflection-based IS. This type of surfaces is also referred to as STAR-RIS as aforementioned [292]. In this subsection, we first introduce the working principles of the omnidirectional IS and its new characteristics. Then, two hardware implementations of the omnidirectional IS are presented. Finally, emerging applications of the omnidirectional IS are discussed.
@@ III-B3 Hardware Implementation
It is worth noting that some existing designs utilize coupling among elements to realize simultaneous reflection and refraction [24]. In other words, the incident signals on an element will also be reflected or refracted from the neighboring elements. As a result, the phase shift matrix $\bm{\Phi}$ is not diagonal any more for either reflection or refraction, and this type of surface is also referred to as BD-RIS as mentioned in Section I-B. Under the BD-RIS framework, the reflective and refractive patches for an element are grouped into a cell, and the cell is modeled as a two-port impedance network, which is similar to the equivalent circuit discussed in [297]. The BD-RIS considers a more general concept that the IS can be divided into several sub-surfaces and each sub-surface is modeled as an $M$-port impedance network, where $M$ is the number of reflective/refractive patches within the sub-surface. This implies that the incident signal on a sub-surface can be reflected or refracted by any patch within each sub-surface. This can be achieved by forming coupling currents among elements within the sub-surface. Such a coupling can provide another dimension to optimize, i.e., the equivalent amplitude of the coefficient for each element can be adjusted, leading to additional performance improvement. However, the coupling poses an increasing complexity on the hardware design of the omnidirectional IS. Therefore, it is necessary to strike a trade-off between the performance and complexity in such a framework.
@@ III-C3 Hardware Implementation
The primary challenge in the hardware implementation of holographic ISs is to design tunable radiation elements which are capable of optimizing the current distribution on the surface. Specifically, the tunability of radiation elements enables the construction of any holographic pattern on the surface, such that dynamic beamforming can be achieved. In the following we will list the state-of-the-art tuning mechanisms for the realization of holographic ISs, which rely on PIN diodes, liquid crystal, and photosensitive devices, respectively. A summary of these hardware implementations is listed in Table VII.
@@
TABLE VII: Comparison Between Different Hardware Implementations
Tuning mechanism
Working frequency
Advantages
Limitation
Typical Applications
PIN diodes
(Sub-) mmWave
Low cost, fast response time and easy implementation
Nonlinearities in high-frequency bands
Mobile communication scenarios
Liquid crystal
(Sub-) mmWave
Low power consumption, linear properties in high-frequency bands
Slow response time
(Quasi-) Static communication scenarios
Photosensitive devices
GHz-optical
Ultra-fast response time, protecting signals from EM interference
Difficult to maintain high phase stability
Long-distance deployments with optical fiber technology
  • PIN diodes: A radiation element capable of controlling the radiation amplitude of the EM wave can be achieved through the integration of a complementary electric-LC (cELC) resonator, which incorporates PIN diodes for modulation purposes [310]. The cELC resonator, as depicted in Fig. 21, is a micro-strip line etched with a circle of annular slot to form a metal patch combined with a closed loop. The operational states of PIN diodes are actively controlled by the application of distinct bias voltages. Such control holds significant influence over the mutual inductance properties of the cELC resonator, consequently dictating the radiation power of the EM wave. Owing to their cost-effectiveness and easy implementation, PIN diodes find extensive applications predominantly in sub-mmWave and mmWave frequency ranges.
  • Liquid crystal: Although PIN diodes are capable of achieving high beampattern gains, the nonlinearities of PIN diodes in high-frequency bands deteriorate the performance of holographic ISs. Thanks to the linear properties of liquid crystal in high-frequency bands, the above limitation of PIN diodes can be overcome. The liquid crystal-based holographic ISs were proposed in [311], where the liquid crystal is filled in cavity boxes between slots and radiating patches. The permittivity of liquid crystal is changed with the applied bias voltages, such that the EM wave can be regulated to achieve beamforming. Liquid crystal-controlled holographic ISs offer numerous benefits, such as their versatility in integration into diverse structures due to their fluid nature and efficiency in power usage, attributing to the predominantly capacitive functions that involve minimal currents. Nonetheless, a significant limitation is the relatively slow response time, which constrains their utility to applications that do not demand rapid tunability.
  • Photosensitive devices: In addition to the above electrical control methods, the tunability of radiation elements can also be realized by optical control. Specifically, different EM responses of each radiation element are controlled by properly illuminating the corresponding photosensitive semiconductors. Such an optical control method is capable of achieving a rapid tuning speed, while also safeguarding signals against EM interference. Moreover, when integrated with optical fiber, it becomes feasible to deploy holographic ISs and signal processing units at separate locations for long-distance deployments. However, a critical challenge associated with these optical tuning systems is the necessity for high phase stability, that is, the preservation of a constant phase difference among parallel optical signals [312]. Since optical fibers are highly susceptible to environmental influences, the above requirement is challenging to be met in practice.
@@ III-C4 Potential Applications
  • LEO satellite communications: LEO satellite communications are one promising solution to deliver high-capacity backhaul or data relay services for terrestrial networks. Nonetheless, the high mobility of LEO satellites and the severe path loss pose challenges on antenna technologies in terms of precise beam steering and high antenna gain [313]. Traditional antennas integrated with user terminals such as dish antennas and phased arrays either require heavy mechanics or costly phase shifters, rendering their adoption in practical systems economically unviable. As a new paradigm to achieve beamforming with low hardware cost and low power consumption, holographic ISs can be adopted in integrated terrestrial-satellite networks to overcome the above limitations. In particular, Kymeta together with Microsoft has developed a commercial holographic IS for satellite communications in the event of natural disasters.
  • Wireless Simultaneous Localization and Mapping (SLAM): Wireless SLAM, as a technique for user positioning and mapping in uncharted environments, holds promise for enhancing location-based services. This method utilizes antennas to estimate both the time of arrival and the angle of arrival by analyzing the amplitudes of received multi-path components [314]. Consequently, the precision of the SLAM system is determined by the directive gain of the antennas. Given that holographic ISs consist of numerous elements, offering enhanced beam-steering capabilities and high directive gain, they present a viable option to be applied in wireless SLAM systems. It also shows great potential in radar detection [315].
  • ISAC: In 6G, it is envisaged that the functions of sensing and communication are fully integrated in the same system by sharing resources, waveforms, hardwares, etc. In ISAC systems, wideband waveforms are required to achieve high data rate for the communication function and high-resolution detection or estimation for the sensing function. However, wideband transmission with traditional phase controlled schemes typically suffers from beam squint. Holographic ISs, with the linear additivity of holographic patterns, can be used to effectively alleviate the beam squint issue [316].